# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### v2

@@ 6 Numerical experiments
Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution. In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 7. We consider $\textbf{h}=1/16$ and other choices for $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ as in Case I. Figure 7 shows that, with the increasing value of $s,$ decreasing trend in the $IT$ counts for both the PESS and LPESS preconditioner. Moreover, for $s>22,$ using LPESS preconditioner, we can solve the three-by-three $SPP$ mentioned only in $3$ iterations.
@@ 7 Conclusion
In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1). For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived. Moreover, we estimated the spectral bounds of the proposed PESS and LPESS preconditioned matrices. This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge. Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS and LPESS preconditioners. The key observations are as follows: $(i)$ the proposed PESS and LPESS preconditioners are found to outperform the existing baseline preconditioners in terms of IT and CPU times. $(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs. $(iii)$ The PESS and LPESS preconditioned matrices have a more clustered spectral distribution than the baseline preconditioned matrices. $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner. Further investigation to determine the optimal parameter $s$ can be done in future research endeavors.

### v3

@@ 7 Numerical experiments
Relationship between $s$ and convergence speed: In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 9. We consider $\textbf{h}=1/16$ and other choices for $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ as in Case I. Figure 9 shows that, with the increasing value of $s,$ decreasing trend in the IT counts for both the PESS and LPESS preconditioner. Moreover, for $s>22,$ using LPESS preconditioner, we can solve this three-by-three block $SPP$ only in $3$ iterations.
Example 3.
[1, 8]
Table 8: Numerical results of GMRES, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS PGMRES processes for Example 3.
Process Problem aug2dc aug3dc listwet12 yao
size($\mathcal{A}$) $50400$ $8746$ $30004$ $6004$
GMRES IT $370$ $78$ $2251$ $49$
CPU $360.1613$ $14.2486$ $422.8209$ $13.9152$
RES $079.9905e-$ 07 $089.2839e-$ 08 $089.9700e-$ 08 $089.7483e-$ 08
IBD IT $5$ $5$ $28$ $25$
CPU $278.7707$ $4.80051$ $278.8961$ $21.9511$
RES $152.1187e-$ 15 $097.5842e-$ 09 $078.3961e-$ 07 $079.4491e-$ 07
MAPSS IT $8$ $8$ $124$ $36$
CPU $324.3828$ $7.6662$ $2078.9607$ $18.6853$
RES $073.6981e-$ 07 $091.7312e-$ 09 $079.2628e-$ 07 $077.6716e-$ 07
SL IT $65$ $14$ $15$ $37$
CPU $2848.2117$ $14.4158$ $270.3901$ $36.7559$
RES $088.4621e-$ 08 $082.0305e-$ 08 $088.2851e-$ 08 $099.8555e-$ 09
SS IT $23$ $6$ $7$ $39$
CPU $1547.7320$ $8.9176$ $200.8205$ $28.9668$
RES $079.9399e-$ 07 $077.7120e-$ 07 $073.5756e-$ 07 $079.5393e-$ 07
RSS IT $23$ $6$ $5$ $16$
CPU $1774.9981$ $8.6199$ $161.8405$ $18.5057$
RES $075.5282e-$ 07 $072.6941e-$ 07 $074.1147e-$ 07 $078.2434e-$ 07
EGSS IT $5$ $5$ $4$ $6$
CPU $350.3081$ $7.61206$ $124.0637$ $6.9551$
RES $123.9934e-$ 12 $121.1711e-$ 12 $071.5556e-$ 07 $072.0227e-$ 07
RPGSS IT $4$ $4$ $3$ $5$
CPU $297.3382$ $6.16249$ $96.5956$ $5.0133$
RES $093.2268e-$ 09 $102.8160e-$ 10 $087.1294e-$ 08 $089.2318e-$ 08
PESS IT $3$ $3$ $2$ $3$
CPU $249.0203$ $5.5716$ $68.5225$ $4.8778$
$s=30$ RES $088.3682e-$ 08 $085.7233e-$ 08 $078.3159e-$ 07 $073.3146e-$ 07
LPESS IT $3$ $3$ $2$ $3$
CPU $237.0945$ $5.4662$ $67.0810$ $3.9688$
$s=30$ RES $091.0780e-$ 09 $098.4676e-$ 09 $074.3037e-$ 07 $073.7858e-$ 07
[…]
within the prescribed IT.
Problem formulation: In this example, we consider the three-by-three block SPP (1.1) arising from the following quadratic programming problem:
$\min_{x\in{\mathbb{R}}^{n},\,y\in{\mathbb{R}}^{p}}\frac{1}{2}x^{T}Ax+f^{T}x+h^{T}y$ (7.2)
such $\text{such that}\leavevmode\nobreak\ \leavevmode\nobreak\ Bx+C^{T}y=b,$ (7.3)
where $f\in{\mathbb{R}}^{n}$ and $h\in{\mathbb{R}}^{p}.$ To solve the minimization problem, we define the Lagrange function:
$L(x,y,\lambda)=\frac{1}{2}x^{T}Ax+f^{T}x+h^{T}y+\lambda^{T}(Bx+C^{T}y-b),$ (7.4)
where $\lambda\in{\mathbb{R}}^{m}.$ Then the Karush-Kuhn-Tucker [47] necessary conditions are as follows:
$\nabla_{x}L(x,y,\lambda)=0,\leavevmode\nobreak\ \nabla_{y}L(x,y,\lambda)=0\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \nabla_{\lambda}L(x,y,\lambda)=0,$
which can be reformulated as a three-by-three block SPP of the form (1.1).
Table 9: Numerical results of PESS-I, LPESS-I, PESS-II and LPESS-II PGMRES processes for Example 3.
Process Problem aug2dc aug3dc listwet12 yao
size($\mathcal{A}$) $50400$ $8746$ $30004$ $6004$
PESS-I IT $5$ $3$ $5$ $6$
($s=1,$ $\Lambda_{1}=0.001I,$ CPU $520.7330$ $10.1460$ $310.5960$ $5.7449$
$\Lambda_{2}=0.1I,$ $\Lambda_{3}=0.001I$) RES $085.4476e-$ 08 $081.1047e-$ 08 $071.5581e-$ 07 $072.9433e-$ 07
LPESS-I IT $4$ $3$ $3$ $3$
($s=1,$ $\Lambda_{2}=0.1I,$ CPU $337.9530$ $8.3194$ $118.2540$ $3.5354$
$\Lambda_{3}=0.001I$) RES $095.4773e-$ 09 $072.8655e-$ 07 $082.0805e-$ 08 $071.5576e-$ 07
PESS-II IT $3$ $3$ $2$ $3$
($s=s_{est},$ $\Lambda_{1}=A,$ CPU $236.4489$ $8.0818$ $74.9301$ $3.7349$
$\Lambda_{2}=\beta_{est}I,$ $\Lambda_{3}=10^{-4}CC^{T}$) RES $083.0982e-$ 08 $095.8697e-$ 09 $097.8956e-$ 09 $076.3723e-$ 07
LPESS-II IT $3$ $3$ $2$ $2$
($s=s_{est},$ $\Lambda_{2}=\beta_{est}I,$ CPU $282.3466$ $8.3194$ $114.7483$ $2.8330$
$\Lambda_{3}=10^{-4}CC^{T}$) RES $093.7409e-$ 09 $091.0390e-$ 09 $081.0465e-$ 08 $086.8900e-$ 08
Data matrices $A$, $B$ and $C$ are selected from the CUTEr collection [48]. The numerical results for the problems ‘aud2dc’, ‘aud3dc’, ‘liswet12’ and ‘yao’ are presented in Table 8.
Parameter selection: The parameter selection for the IBD and MAPSS preconditioners are as in Example 1. For the SS, RSS, EGSS, RPGSS, PESS and LPESS preconditioners, the parameters are chosen the same as in Case II of Example 1 with $\alpha=0.1$ and $\Lambda_{1}=0.1I$.
Results for experimentally found optimal parameter: For the proposed preconditioner, the experimentally found parameter $s$ is $30$ in the interval $[10,30].$ The numerical results presented in Table 8 demonstrate that the proposed PESS and LPESS preconditioners significantly outperform existing baseline preconditioners in terms of IT and CPU time. The proposed preconditioners can solve the there-by-three block SPP within $2$ or $3$ iterations. Furthermore, for the problem ‘liswet12’, our proposed LPESS preconditioner is approximately $76\%$, $97\%$, $75\%,$ $66\%,$ $59\%,$ $46\%$ and $31\%$ more efficient than the existing IBD, MAPSS, SL, SS, RSS, EGSS and RPGSS preconditioner, respectively.
Results using parameters selection strategy in Section 6: To further illustrate the effectiveness of the proposed preconditioners under the parameter selection strategy discussed in Section 6, we reported the numerical results in Table 9 for the preconditioners PESS-I, LPESS-I, PESS-II and LPESS-II (as discussed in Example $1$). A comparison of the results in Table 9 with the existing baseline preconditioners listed in Table 8 indicates that the proposed preconditioners perform well when the parameter $s$ is selected as outlined in Section 6.
@@ 8 Conclusion
In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1). For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived. Moreover, we estimated the spectral bounds of the proposed PESS and LPESS preconditioned matrices. This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge. Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS and LPESS preconditioners. The key observations are as follows: $(i)$ the proposed PESS and LPESS preconditioners are found to outperform the existing baseline preconditioners in terms of IT and CPU times. $(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing their proficiency in solving three-by-three block SPPs. $(iii)$ The proposed PESS and LPESS preconditioned matrices have better clustered spectral distribution than the baseline preconditioned matrices. $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner.